OpenStudy (anonymous):

Find an equation of the lines passing thru each pair of points. Write eqution in form Ax+Bx=C... Here are the points (2,3) and (-1,-1)

OpenStudy (anonymous):

have you heard on the wine glasses on a pinic table? its pretty much Y2-Y1/X2-X1

OpenStudy (anonymous):

lol, when i do the problem it give me -2x+y=-1, and in the back of the book its 4x-3y=-1

OpenStudy (radar):

The book knows. solve for slope as RedCarr posted. Then 4/3=(y-3)/(x-2)